OpenStudy (poofypenguin):

Find conditions on the variable “k” under which the intersection of the given set of planes is: a) a point b) a line c) does not exist Plane 1: kx – y +z = -3 Plane 2: x + y – kz= 1 Plane 3 : 3x + 2y – 2z = 0

OpenStudy (anonymous):

b) determinant \[\left[\begin{matrix}k &-1&1 \\ 1 & 1&-k \\ 3&2&-2\end{matrix}\right]=0\] It happens for k=1 and k=-3/2

OpenStudy (poofypenguin):

ummm... i'm not quite sure what your solution means?

OpenStudy (anonymous):

if the planes intersect in a line, it means that their perpendicular vectors all belong to one plane. So it means this vectors are dependent. That's why determinant formed with them have to be equal 0

OpenStudy (poofypenguin):

i'm unframiliar with the term determinant... my school does not include matix in the curiculum...

OpenStudy (anonymous):

then look here, maybe helps... http://answers.yahoo.com/question/index?qid=20080825150238AAe2UdK

OpenStudy (poofypenguin):

i've got those basics already... i'm just not sure how to solve for k